OpenStudy (anonymous):

can someone help me find the solution? 4(x+1)-3=7(x-1) i came out with -5 =x is this right?

OpenStudy (kkutie7):

\[4(x+1)-3=7(x-1)\] \[4x+4-3=7x-7\] \[4x+1=7x-7\] \[8=3x\] \[\frac{8}{3}=x\] this is what I got, but I'm tired so I'd check it first.

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

tired or not, it is right

OpenStudy (kkutie7):

your welcome